• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 07:58:25 on July14,2001

    Entry number 66104

    keyword=shift summary

    00:00 IOC problem. Beam down.
    00:35 Beam restored to hall
    01:35 NMR measurement:49.2%
    01:45 Beam restored to hall
    02:10 Beam down for Compton Daq reboot.
    02:15 Beam restored to hall
    02:25 Problems with IOC again. Beam down
    03:00 Beam restored to hall
    04:00 EPR:48.8%
    04:15 Large helicity asymmetry measured.
    Began series of short runs to reduce asymm.
    Waveplate frozen at 750.
    Called MCC to check status of waveplate server.
    04:35 Problems with IOC. Beam Down
    04:40 Beam back.
    04:57 MCC reboots the lock server.
    05:15 MCC attempting to modify waveplate by hand
    05:40 MCC completes manual waveplate rotation.
    Asymmetry much improved.
    Disabled helicity feedback.
    06:45 NMR: 46.6%
    06:50 Beam restored to hall


    NOTE: Helicity feedback software has been disabled.
    It should be reacitvated when communication
    with rotatable waveplate is re-established.
    Until then use `apar' to monitor asymm. MCC can perform
    manual adjustment for large asymmetries.